The equation h=-16t^2+112t gives the height of an arrow, shot upward from the groung with an initial velocity of 112 ft/s, where t is the time after the arrow leaves the ground. Find the time it takes for the arrow to reach a height of 180 ft.
Answer by checkley75(3666)

180=-16T^2+112T
-16T^2+112T-180=0
(-8T+20)(2T-9)=0
-8T+20=0
-8T=-20
T=-20/-8
T=2.5 ANSWER.
2T-9=0
2T=9
T=9/2
T=4.5 ANSWER.
-------------------------------------------------------
2.5 seconds on the way up & 4.5 seconds on the way down the arrow will be 180 feet above ground.